OpenStudy (anonymous):

can somebody please help me with a few math questions? i suck at it :( I WILL GIVE A MEDAL AND FAN!!!!! 1.is the equation true,false,or open 4y+8=6y+3 a.true:the expressions are the same for all values and variables. b.false;the expressions are never the same c.open;there is a variable i think its open 2.which value is a solution of theequation5-4x=-3 a.0 b.2 c.-3 d.1/4 3.which ordered pair is a solution of the equation y=x-4 a.(2,6) b.(6,2) c.(-2,6) d(-6,2)4.which oredered pair is a solution of the equation y=5x a.(-2,10) b.(-5,25) c.(-3,15) d.(-2,-10) please explain how you got your answer.

ganeshie8 (ganeshie8):

you're correct for #1

ganeshie8 (ganeshie8):

for #2 : 5 - 4x = -3

ganeshie8 (ganeshie8):

when they ask u for solution, they want u to isolate the "variable"

ganeshie8 (ganeshie8):

5 - 4x = -3 start by subtracting 5 both sides : 5 - 4x = -3 -5 -5 -4x = -8

OpenStudy (anonymous):

thank you!! @ganeshie8

ganeshie8 (ganeshie8):

np :) you're able to finish them ha ?